OpenStudy (jkristia):

Problem set 1, Question 1F-7 d. I'm not able to get the correct result. I get (-6a - 6at) in the numerator of the first fraction, where the result is (at * -6bt) Isn't it correct that the product rule for 3 products is f(x) = uvw f'(x) = u v' w' + u' v w' + u' v' w.

OpenStudy (jkristia):

The equation is \[at(1+bt^2)^{-3}\] I get the derivative for the last product \[w' = [-3(1+bt^2)^{-4}(2) = -6(1+bt^2)^{-4} \] \[a(1)w' + atw' + \frac{a(1)}{(1+bt^2)^3}\] so I get \[\frac{-6a}{(1+bt^2)^4}+\frac{-6at}{(1+bt^2)^4}+\frac{a}{(1+bt^2)^3}\] What am I missing?

OpenStudy (jkristia):

ahhh, see - sometimes it help to ask. I think I found my mistake. 'a' is of course just a constant. Let me try one more time

OpenStudy (jkristia):

so, got the chain rule wrong (again!! :) ) \[w' = -3(1+bt^2)^{-4}(2bt) = \frac{-6bt}{(1+bt^2)^4}\] And now I get the correct result.